OpenStudy (anonymous):

HELP PLEASE! Which of the following is NOT true concerning the equation x^2-c=0 for c>0? A. The left-hand side of this equation is called a difference of two squares. B. A quadratic equation in this form can always be solved by factoring. C. A quadratic equation in this form can always be solved using the square root property. D. This equation is not considered to be a quadratic equation because it is not of the form zx^2+bx+c=0. (I have a feeling about D, but I don't know, might be utterly wrong.)

jimthompson5910 (jim_thompson5910):

It's not D because x^2 - c is the same as 1x^2 + 0x + (-c) notice how it matches the form ax^2 + bx + c where a = 1, b = 0, c is really -c

jimthompson5910 (jim_thompson5910):

so x^2 - c = 0 is actually a quadratic equation

OpenStudy (anonymous):

Oh yeah thought so wrote that down , thanks for the correction so it either be B or C?

jimthompson5910 (jim_thompson5910):

x^2 - c is only factorable if c is a perfect square example: x^2 - 4 = (x-2)(x+2) in any other case, it's not factorable The good news is that x^2 - c = 0 can be always be solved using the square root property

jimthompson5910 (jim_thompson5910):

Like so.. \[\Large x^2 - c = 0\] \[\Large x^2 - c+c = 0+c\] \[\Large x^2 = c\] \[\Large \sqrt{x^2} = \pm\sqrt{c}\] \[\Large x = \pm\sqrt{c}\]

jimthompson5910 (jim_thompson5910):

all that shows the square root property in action